OpenStudy (garretttyre05):

Need help asap.@faiqraees What is p−^m ÷ p^n equal to? A. p−^m + ^n B. p−^m ⋅ ^n C. p−^m − ^n D. p−^m ÷ ^n

OpenStudy (garretttyre05):

@FaiqRaees

OpenStudy (faiqraees):

are you sure you have given us the correct question?

OpenStudy (garretttyre05):

there i edited it

OpenStudy (faiqraees):

is it p-^m or p^-m

OpenStudy (garretttyre05):

\[p^{-m} \div p^{n}\]

OpenStudy (faiqraees):

\[\large\rm a^x / a^y = a^{x-y} \] Use the formula

OpenStudy (garretttyre05):

so B?

OpenStudy (garretttyre05):

because you would change it to multiplication

OpenStudy (garretttyre05):

@FaiqRaees

OpenStudy (faiqraees):

No not b.

OpenStudy (faiqraees):

p^−m÷p^n = p^(-m-....?)

OpenStudy (garretttyre05):

im confused

OpenStudy (faiqraees):

lets suppose x = -m and y = n Now substitute these values in the equation I gave

OpenStudy (garretttyre05):

\[p ^{m} \div p ^{n}=p ^{m-n}\]

OpenStudy (faiqraees):

substitute in the equation I gave you. The one which had a's in it

OpenStudy (garretttyre05):

\[A ^{M} \div A ^{N}=A ^{M-N}\]

OpenStudy (garretttyre05):

So the answer is C

OpenStudy (faiqraees):

p^−m÷p^n a^x ÷ a^y = a^(x-y) x = -m y = n a^(-m) ÷ a^n = a^((-m)-n)

OpenStudy (garretttyre05):

C? @FaiqRaees

OpenStudy (faiqraees):

can you please mention the options more correctly

OpenStudy (garretttyre05):

OK @FaiqRaees